You are here

Beam Delivery News

Primary tabs

TItle Date Excerpt Attached images
Beam delivery update: Maintenance tomorrow 02 September 2014

The cyclotron is currently delivering nA as needed to BL1B for PIF, 9.8 µA to BL2A for RIB production, and ~95 µA to BL2C4 for isotope production at STF. This is expected to continue until ~06:00 tomorrow, at which point BL2A will be defined off for access to the mass separator room and a target module move. BL1A and BL2C4 will continue to run until ~08:30.

Beam delivery update 31 August 2014

The cyclotron is delivering ~105 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4, while at ISAC 31Na is being delivered to Osaka experiment S1391. Where we have a problem is with the ISAC-I RFQ.

Beam delivery update (SPECIAL LONG WEEKEND EDITION) 29 August 2014

The cyclotron is running normally, delivering ~105 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4. Other than the possibility of an STF target change at some point (which would see BL2C4 defined off and might cause brief disruptions on the other beamlines), this is expected to continue through the weekend.

Beam delivery update 28 August 2014

The cyclotron is currently delivering ~110 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4 for isotope production at STF. With the exception of an STF target change at some point, this is expected to continue through the upcoming long weekend.

Beam delivery update: beam development completed 26 August 2014

Today’s beam development was finished shortly before 20:00 and full operation is being restored. The plan is to deliver ~110 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4 as conditions allow.

Beam delivery update: Maintenance during beam development tomorrow 25 August 2014

The cyclotron’s running normally, delivering ~110 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4, while at ISAC 34Al is being delivered to TITAN for experiment S1240. These are both scheduled to continue until the start of tomorrow’s beam development shift at 08:30. There is no plan to interrupt operation for Wednesday’s “maintenance if required”.

Beam delivery update 21 August 2014

The cyclotron is running normally, delivering ~110 µA to BL1A, 9.8 µA to BL2A, and ~98 µA to BL2C4. At ISAC, the new UCtarget in TM4 saw its first protons last night. Both cyclotron operation and RIB delivery are scheduled to continue through the weekend.

Beam delivery update: Maintenance tomorrow 19 August 2014

The cyclotron is delivering ~110 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4. This is scheduled to continue until 08:30 tomorrow, at which point the cyclotron will go off for an ISIS backplate change. 

Beam delivery update: weekend edition 15 August 2014

The cyclotron is currently delivering ~110 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4. Ops is working to address a low water flow issue.

Beam delivery update: operation restored 13 August 2014

The cyclotron is back to full operation and delivering ~110 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4. We’re scheduled to run like this until next week’s maintenance.

Pages

Subscribe to Group News